OpenStudy (anonymous):

WILL FAN AND MEDAL 1 QUESTION! What is the value of -(-2)3<---thats a little three hanging off the top, like 3 squared i believe.

OpenStudy (rrrrrrrrr):

-6

OpenStudy (rrrrrrrrr):

-8 sorry

OpenStudy (anonymous):

\(-(-2)^3\) ?

OpenStudy (anonymous):

The exponent of 3 means "cubed". Multiply -2 by itself 3 times. Don't forget the negative on the outside, which is understood as -1. \[-(-2)^3=-1\times-2\times-2\times-2=?\]

OpenStudy (anonymous):

@rrrrrrrrr that was wrong:( thanks for trying tho:)